40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Fri May 26 21:53:20 2017, Koji, General, Configuration, Trans RF PD setup 
    Reply  Tue Jun 6 00:49:48 2017, Koji, General, Configuration, Trans RF PD setup 
       Reply  Tue Jun 6 22:00:36 2017, Koji, General, Configuration, Trans RF PD setup 
Message ID: 278     Entry time: Fri May 26 21:53:20 2017     Reply to this: 279
Author: Koji 
Type: General 
Category: Configuration 
Subject: Trans RF PD setup 

Recent work

- DC output of the trans RF PD was connected to the BNC patch panel. => Now CH4 of the scope is monitoring this signal

- The RF sweep signal from the network analyzer is connected to the power combiner for the EOM drive via the SMA patch panel.

- The trans RF PD was aligned first to the leakage beam. It turned out that this signal is too weak. Then the PD was aligned to one of the main OMC transmission. For this purpose, the OMC DCPD (T) was removed from the OMC breadboard.

- It seems that there is a significant amount of RF AM from the EOM. I suspect it is associated with the residual S-pol and birefringence of the steering mirrors (45deg HR). But the HWP at the output of the Faraday is fixed on the Faraday body with a screw and cumbersome for fine adjustment. A PBS and an HWP are added right before the EOM. This made the fiber coupler slightly misaligned. I suppose this new setup still has S&P on the fiber too. Thus, readjustment of the fiber rotations at the input is necessary.

Next step

- Input power to the fiber should be determined before the EOM. Otherwise, touching the HWP before the EOM causes too much power change at the optics of the OMC side.

- Precise adjustment of the RFAM is still necessary.

- The OMC curved mirror should be held by the new fixture.

- Check the beam spots

- Measure cavity parameters. (transmission/FSR/HOM/etc)

==> Then the curved mirror and the PZT will be glued on the prism

ELOG V3.1.3-